# Edge-of-chaos enhanced quantum-inspired algorithm for combinatorial optimization

### v4

@@ Abstract
Nonlinear dynamical systems with continuous variables can be used for solving combinatorial optimization problems with discrete variables. Numerical simulations of them are also useful as heuristic algorithms with a desirable property, namely, parallelizability, which allows us to execute them in a massively parallel manner, leading to ultrafast performance. However, the dynamical-system approaches with continuous variables are usually less accurate than conventional approaches with discrete variables such as simulated annealing. To improve the solution accuracy of a quantum-inspired algorithm called simulated bifurcation (SB), which was found from classical simulation of a quantum nonlinear oscillator network exhibiting quantum bifurcation, here we generalize it by introducing nonlinear control of individual bifurcation parameters and show that the generalized SB (GSB) can achieve surprisingly high performance, namely, almost 100% success probabilities for some large-scale problems. As a result, the time to solution for a 2,000-variable problem is shortened to 10 ms by a GSB-based machine, which is two orders of magnitude shorter than the best known value, 1.3 s, previously obtained by an SB-based machine. To examine the reason for the ultrahigh performance, we investigated chaos in the GSB changing the nonlinear-control strength and found that the dramatic increase of success probabilities happens near the edge of chaos. That is, the GSB can find a solution with high probability by harnessing the edge of chaos. This finding suggests that dynamical-system approaches to combinatorial optimization will be enhanced by harnessing the edge of chaos, opening a broad possibility to tackle intractable combinatorial optimization problems by physics-inspired approaches.
@@ V No influence of discretization
Figures 1(g–i) also show that $P_{S}$ does not monotonically increase with respect to the number of steps $M$, unlike SA. To investigate the reason for this, we checked the normalized distance at the final time $\delta(t_{M})$ corresponding to Figs. 1(g–i), the results of which are shown in Figs. 1(j–l). Interestingly, Fig. 1(j) suggests that $P_{S}$ increases around the “valley of chaos.” This is also the case in Figs. 1(k) and 1(l), though they are less clear. This is natural in the sense that the butterfly effect becomes more evident for longer time evolution. Thus, a too large $M$ (more precisely $t_{M}$) may result in chaos and consequently $P_{S}$ may become lower.
@@ VII Summary and outlook
Our results open broad and interesting possibilities as follows. In this work, we have focused on bSB and found the edge-of-chaos effect. It is thus interesting whether or not other dynamical-system approaches, such as aSB [22], dSB [24], neural networks [4, 5], coherent Ising machines [6, 7, 8, 9, 10, 11], and oscillator-based Ising machines [33, 34], exhibit similar effects by similar nonlinear control. The GbSB has an additional parameter, namely, the nonlinear-control strength, which requires additional tuning for achieving high performance, together with tuning of the number of time steps. Although we have not included the time to optimize these parameters in computation times in this work, as in previous studies, the optimization time is not negligible for practical applications. To optimize the parameters quickly, our finding that good parameter values may be found around the edge of chaos will be helpful. The GbSB could not improve G9, one of the G-set instances, compared with dSB. This result suggests the limitation of the power of GbSB. While it is important to clarify the range of applicability of GbSB, this result suggests that we may need a hybrid algorithm combining GbSB with others for solving a wide range of problems. In general, the SBs can achieve high performance for dense-connectivity problems [3, 69]. On the other hand, a dynamical-system approach that can achieve ultrahigh performance for sparse-connectivity problems has recently been reported [70]. The combination of them will be promising to tackle both kinds of problems.
@@ Appendix A Parameter settings in the GbSB
As for initial conditions, $x_{i}(0)$ was set to a random value between 1-1 and 1, except the simulations to evaluate $\delta(t_{M})$ in Figs. 1(d–f), and $y_{i}(0)$ was set to zero.

### v5

@@ Abstract
Nonlinear dynamical systems with continuous variables can be used for solving combinatorial optimization problems with discrete variables. Numerical simulations of them are also useful as heuristic algorithms with a desirable property, namely, parallelizability, which allows us to execute them in a massively parallel manner, leading to ultrafast performance. However, the dynamical-system approaches with continuous variables are usually less accurate than conventional approaches with discrete variables such as simulated annealing. To improve the solution accuracy of a quantum-inspired algorithm called simulated bifurcation (SB), which was found from classical simulation of a quantum nonlinear oscillator network exhibiting quantum bifurcation, here we generalize it by introducing nonlinear control of individual bifurcation parameters and show that the generalized SB (GSB) can achieve surprisingly high performance, namely, almost 100% success probabilities for some large-scale problems. As a result, the time to solution for a 2,000-variable problem is shortened to 10 ms by a GSB-based machine, which is two orders of magnitude shorter than the best known value, 1.3 s, previously obtained by an SB-based machine. To examine the reason for the ultrahigh performance, we investigated chaos in the GSB changing the nonlinear-control strength and found that the dramatic increase of success probabilities happens near the edge of chaos. That is, the GSB can find a solution with high probability by harnessing the edge of chaos. This finding suggests that dynamical-system approaches to combinatorial optimization will be enhanced by harnessing the edge of chaos, opening a broad possibility for physics-inspired approaches to combinatorial optimization.
@@ V No influence of discretization
Figures 1(g–i) also show that $P_{S}$ does not monotonically increase with respect to the number of steps $M$, unlike SA. To investigate the reason for this, we checked the normalized distance at the final time $\delta(t_{M})$ corresponding to Figs. 1(g–i), the results of which are shown in Figs. 1(j–l). Interestingly, Fig. 1(j) suggests that $P_{S}$ increases around the “valley of chaos.” This is also the case in Figs. 1(k) and 1(l), though they are less clear. Note that a longer $t_{M}$ (larger $M$) leads to lower decreasing rates of the bifurcation parameters through Eq. (7). This change of the decreasing rates may affect the chaos in the present system and lead to the valley of chaos. So far, however, it is difficult to understand how the success probability $P_{S}$ increases through this change, which is left for future work.
@@ VII Summary and outlook
Our results open broad and interesting possibilities as follows. In this work, we have focused on bSB and found the edge-of-chaos effect. It is thus interesting whether or not other dynamical-system approaches, such as aSB [22], dSB [24], neural networks [4, 5], coherent Ising machines [6, 7, 8, 9, 10, 11], and oscillator-based Ising machines [33, 34], exhibit similar effects by similar nonlinear control. The GbSB has an additional parameter, namely, the nonlinear-control strength, which requires additional tuning for achieving high performance, together with tuning of the number of time steps. Although we have not included the time to optimize these parameters in computation times in this work, as in previous studies, the optimization time is not negligible for practical applications. So far, the parameter optimization relies on grid search, as shown in Figs. 1(a–c). To optimize the parameters more quickly, our finding that good parameter values may be found around the edge of chaos will be helpful. The GbSB could not improve G9, one of the G-set instances, compared with dSB. This result suggests the limitation of the power of GbSB. While it is important to clarify the range of applicability of GbSB, this result suggests that we may need a hybrid algorithm combining GbSB with others for solving a wide range of problems. In general, the SBs can achieve high performance for dense-connectivity problems [3, 69]. On the other hand, a dynamical-system approach that can achieve ultrahigh performance for sparse-connectivity problems has recently been reported [70]. The combination of them will be promising to tackle both kinds of problems.
@@ Appendix A Parameter settings in the GbSB
As for initial conditions, $x_{i}(0)$ was set to a random value between 0.1-0.1 and 0.1, except the simulations to evaluate $\delta(t_{M})$ in Figs. 1(d–f), and $y_{i}(0)$ was set to zero.